# Searcher Competition in Block Building

### v1

@@ Abstract
We study the amount of maximal extractable value (MEV) captured by validators, as a function of searcher competition, in blockchains with competitive block building markets such as Ethereum. We argue that the core is a suitable solution concept in this context that makes robust predictions that are independent of implementation details or specific mechanisms chosen. We characterize how much value validators extract in the core and quantify the surplus share of validators as a function of searcher competition. Searchers can obtain at most the marginal value increase of the winning block relative to the best block that can be built without their bundles. Dually this gives a lower bound on the value extracted by the validator. If arbitrages are easy to find and many searchers find similar bundles, the validator gets paid all value almost surely, while searchers can capture most value if there is little searcher competition per arbitrage. For the case of passive block-proposers we study, moreover, mechanisms that implement core allocations in dominant strategies and find that for submodular value, there is a unique dominant-strategy incentive compatible core-selecting mechanism that gives each searcher exactly their marginal value contribution to the winning block.
@@ 1 Introduction
Blockchains that support smart contracts frequently run decentralized finance (DeFi) applications. This in turn gives rise to the phenomenon of miner/maximal extractable value, [5]: Blockchain protocols give validators, sometimes called proposers, the right to order transactions for a particular block. This effectively means that validators have a local monopoly to include or exclude transactions or to order transactions in a particular way, in order to generate value for themselves from this privileged position. However, since extracting value from transaction ordering in an optimal way is a difficult task, more specialized actors such as block builders and (arbitrage) searchers participate in the value extraction process in smart contract blockchains such as Ethereum. Block builders aggregate different arbitrage opportunities, liquidations or ”sandwiches” in one block, using transactions from the public mempool, different private mempools, order flow auctions or other transaction sources together with their own transactions. Then, they bid against other builders, to get their block published in a canonical chain of blocks. The bid is paid to the current block proposer. Arbitrage opportunities are typically found by more specialized players, searchers and passed to the builders. See Figure 1.
@@
In this paper, we analyze the competition between searchers in order to explain the value distribution between the proposer (or block builder) and the searchers. Alternatively, our model can also be applied to the competition of different order flow providers and explain the value distribution between different order flow providers and the proposer (or block builder). In our model, different searchers may capture different value from different opportunities and may or may not find unique opportunities that are not found by competing searchers. This competition and specialization should in turn explain the value capture by different searchers in the MEV supply chain. Our model abstracts away from the intermediate layer of block building where block builders aggregate searcher bundles to blocks and compete in a bidding procedure. However, the ability for different combination of searchers or order flow providers to jointly generate different blocks is captured implicitly by the solution concept we focus on: We consider core allocation where different combinations of players can construct blocks of different values and distribute the value among themselves. The core constraints require that no coalition of players should be able to generate more value for themselves (produce a more valuable block) than they get paid in the realized allocation. Thus, while we generically talk about searchers, our model can for example also capture the case where some searchers act as block builders. The core is a natural solution concept in the context of block building and MEV capture as it captures the competition and possibility of collaboration of different players while abstracting away from any particular mechanisms that would intermediates between searchers, builders and proposers. Thus, it provides robust predictions on the set of plausible value distributions that can arise in any sufficiently competitive block building market. [^2] The ability to make predictions about value distributions could be particularly useful in situations where profits are not directly observed. It has for example been observed that block builders with access to exclusive order flow tend capture more value in the block building market. However, it cannot necessarily be inferred from on chain data how much of the profit the builder actually passes on to the order flow provider, since side payments might not be directly observable. Our model would generally predict that order flow provider with unique flow should capture most of the value they contribute. As a next step, one can then look into mechanism that implement particular core allocations through elicitating privately held information of searchers about the value they can generate from different blocks.
@@
Our first result, shows that searchers can at most obtain their marginal contribution to the winning block (the difference in value between the best block that can be built with their transactions and without their transactions) and if value is submodular, then giving each searcher their marginal contribution is in the core. [^3] While transactions submitted to a builder might exhibit significant complementarities, submodularity can be argued to be a reasonable assumption in a sufficiently consolidated market. Order flow providers with complementary flow have an incentive to integrate to capture more value together. Thus, the realizable value might be submodular in the contributions of the different (consolidated) players. A straightforward calculation shows that in a world with passive block producers, this particular core allocation coincides with the Vickrey-Clark-Growes (VCG) outcome ([13]) and hence can be implemented in dominant strategies (Corollary 1). On the other hand (Proposition 2), no other core-selecting mechanism is dominant-strategy incentive compatible. These result gives additional justification to further study the searcher-optimal point in the core: the value obtained by the validator in this point can be interpreted as the maximal extractable value he can obtain taking into account information rents captured by the searchers.
@@
As a next step, we consider a stochastic model of searcher competition to study the value distribution between the searchers and the validator. Different (arbitrage or MEV) opportunities have a fixed probability of being found by a searcher. The more searchers find an opportunity, the more value can be extracted by the validator and the less value can be extracted by the searchers. We show that if the probability of finding an opportunity is bounded from below by $p>\log(n)/n$, where $n$ is the number of searchers actively searching in the strategy, with high probability the validator captures all value in all core allocations. On the other hand, if the probability of sucessfully searching is very low, $p\in\Theta(1/n)$, then with positive constant probability, searchers will be able to capture all value in the searcher-optimal core allocation.
@@
Finally, we validate our stochastic model empirically. We obtain bundle data from the MEV-Share Order Flow Auction (OFA) that allows to study the relation between searcher competition and validator profit. We find that there is significant positive relation between the number of bundles that backrun the same user transaction and the median profit of the validator from opportunities with the same number of backruns.
@@ 1.1 Related Literature
The phenomenon of miner extractable value has first been documented in [5]. Another early contribution on front-running in decentralized finance is [8]. MEV has been documented by a variety of public dashboards and data sets, see for example https://mevboost.pics/,https://https://libmev.com/. Recently, also the magnitude of non-atomic arbitrage has been empirically investigated, where searchers realize one lag of a trade on chain and one on a different domain, see e.g.[10] and https://dune.com/flashbots/lol.
The block building market structure that has evolved in Ethereum since the change to proof of stake, has been the topic of several recent contributions: The dashboard https://orderflow.art/ documents empirically the supply chain through which transaction requests land on chain. [1] studies whether MEV and Proof-of-Stake rewards capture leads to centralization, discussing both validator and builder roles in the market formation. [14] argue that in the current market structure searchers and builders have an incentive to vertically integrate. [4] argue that the builder market is prone to centralization. [17] provide descriptive statistics on the level of decentralization on the builder landscape. [3] proposes a dynamic MEV sharing mechanism that the authors argue results into better decentralization and fair allocation.
@@
The notion of the core was first formally defined in [9]. For an overview of results around the core and submodularity see e.g. [11]. Similar considerations as in our paper, appear in the literature on core-selecting auctions [7, 6].
2 Model
@@
i.e. in case the validator is part of the coalition, the coalitional value is the value of the total welfare maximizing block consisting of transactions from searchers in the coalition, and in case the validator is not part of the coalition, no value can be generated, as the validator is necessary to realize a block. It will be useful subsequently to introduce the short hand notation
@@
if we receive more bundles to build a block that will increase welfare weakly, since we can always discard submitted bundles when building a block. Moreover, we make the following assumption on the (collective) value function:
@@
Submodularity states that the value of a block we can build from transactions from searchers in $A$ and $B$, is bounded by subtracting the value of a block we can build from transactions in both $A$ and $B$ from the sum of values we can achieve from building a block with transactions in $A$ and a block with transactions in $B$.
Submodularity requires that there are not-too-strong complementarities between different submitted bundles. We can justify this assumption in two ways: first, it may be that complementarities are not strong and different MEV opportunities provide value that is mostly independent from other opportunities. Second, it may be that the complementarities are already absorbed by searchers, e.g. in the sense that searchers who provide complementary flow have an incentive to integrate their operations and send their flow together to extract more value. It is worth noting that our upper bound on searcher values holds also for non-submodular value functions, but may be loose in that case.
@@
Decreasing Marginal Value: Let $A\subseteq B\subseteq\mathcal{S}$ and $a\in A$. Then
@@
A direct consequence of submodularity is the following lemma which will be useful subsequently:
@@
Let ${A}\subseteq{B}\subseteq\mathcal{S}$. Then
@@
We prove the result by induction on $N:=|{B}\setminus{A}|$. For $N=0$ the result holds trivially. Now suppose the result hold for $N\geq 0$ and consider the case $N+1$. Let $j\in B\setminus A$. By induction assumption
@@ 3 Analysis
Our first main results states each searcher can at most capture their marginal contribution to the realized block, and any allocation that gives each searcher at most their marginal contribution is in the core. In particular, this implies that there is a searcher optimal core allocation (giving each searcher exactly their marginal contribution to the realized block) and a validator optimal allocation (giving the validator all realized value).
@@
First we show that $x_{i}\leq\bar{v}(\mathcal{S})-\bar{v}(\mathcal{S}\setminus\{i\})$ for each $i$. Suppose not. In that case
@@
where the last inequality follows from Lemma 1. Thus, the core inequalities (1) are satisfied for all coalitions $A\cup V$ with $A\subseteq\mathcal{S}$. For coalitions without the validator the core inequalities are trivially satisfied, as all agents get non-negative value in $x$.

### v2

# Akaki Mamageishvili Offchain Labs Christoph Schlegel Flashbots Benny Sudakov ETH Zürich Danning Sui Flashbots

@@ Abstract
We study the amount of maximal extractable value (MEV) captured by validators, as a function of searcher competition, in blockchains with competitive block building markets such as Ethereum. We argue that the core is a suitable solution concept in this context that makes robust predictions that are independent of implementation details or specific mechanisms chosen. We characterize how much value validators extract in the core and quantify the surplus share of validators as a function of searcher competition. Searchers can obtain at most the marginal value increase of the winning block relative to the best block that can be built without their bundles. Dually this gives a lower bound on the value extracted by the validator. If arbitrages are easy to find and many searchers find similar bundles, the validator gets paid all value almost surely, while searchers can capture most value if there is little searcher competition per arbitrage. Moreover, mechanisms that implement core allocations in dominant strategies, for submodular values, there is a unique dominant-strategy incentive compatible core-selecting mechanism that gives each searcher exactly their marginal value contribution to the winning block. We extend our model to multiple concurrent proposers in which, under mild assumptions, the core is empty.
@@ 1 Introduction
Blockchains that support smart contracts frequently run decentralized finance (DeFi) applications. This in turn gives rise to the phenomenon of miner/maximal extractable value, [5]: Blockchain protocols give validators, sometimes called proposers, the right to order transactions for a particular block. This effectively means that validators have a local monopoly on including or excluding transactions or ordering transactions in a particular way, in order to generate value for themselves from this privileged position. However, since extracting value from transaction ordering in an optimal way is a difficult task, more specialized actors, such as block builders and (arbitrage) searchers participate in the value extraction process in smart contract blockchains such as Ethereum. Block builders aggregate different arbitrage opportunities, liquidations or ”sandwiches” in one block, using transactions from the public mempool, different private mempools, order flow auctions, or other transaction sources together with their own transactions. Then, they bid against other builders, to get their block published in a canonical chain of blocks. The bid is paid to the current block proposer. Arbitrage opportunities are typically found by more specialized players, searchers and passed to the builders. See Figure 1.
@@
In this paper, we analyze the competition between searchers in order to explain the value distribution between the proposer (or block builder) and the searchers. Alternatively, our model can also be applied to the competition of different order flow providers and explain the value distribution between different order flow providers and the proposer (or block builder). In our model, different searchers may capture different value from different opportunities and may or may not find unique opportunities that are not found by competing searchers. This competition and specialization should, in turn, explain the value capture by different searchers in the MEV supply chain. Our model abstracts away from the intermediate layer of block building where block builders aggregate searcher bundles to blocks and compete in a bidding procedure. However, the ability for different combinations of searchers or order flow providers to jointly generate different blocks is captured implicitly by the solution concept we focus on: We consider core allocation where different combinations of players can construct blocks of different values and distribute the value among themselves. The core constraints require that no coalition of players should be able to generate more value for themselves (produce a more valuable block) than they get paid in the realized allocation. Thus, while we generically talk about searchers, our model can, for example, also capture the case where some searchers act as block builders. The core is a natural solution concept in the context of block building and MEV capture as it captures the competition and possibility of collaboration of different players while abstracting away from any particular mechanisms that would intermediate between searchers, builders and proposers. Thus, it provides robust predictions on the set of plausible value distributions that can arise in any sufficiently competitive block building market. [^2] The ability to make predictions about value distributions could be particularly useful in situations where profits are not directly observed. It has, for example, been observed that block builders with access to exclusive order flow tend to capture more value in the block building market. However, it cannot necessarily be inferred from on chain data how much of the profit the builder actually passes on to the order flow provider, since side payments might not be directly observable. Our model would generally predict that an order flow provider with unique flow should capture most of the value they contribute. As a next step, one can look into mechanisms that implement particular core allocations by eliciting privately held information of searchers about the value they can generate from different blocks. All values in the model are assumed to be calculated after paying chain (gas) fees. One extra way to have heterogeneous values across searchers is their ability to reduce (gas) fees they pay.
@@
Our first result shows that searchers can at most obtain their marginal contribution to the winning block (the difference in value between the best block that can be built with their transactions and without their transactions). If the value is submodular (value in the set extension has diminishing returns), then giving each searcher their marginal contribution is in the core. [^3] While transactions submitted to a builder might exhibit significant complementarities, submodularity can be argued to be a reasonable assumption in a sufficiently consolidated market. Order flow providers with complementary flow have an incentive to integrate to capture more value together. Thus, the realizable value might be submodular in the contributions of the different (consolidated) players. A straightforward calculation shows that in a world with passive block producers, this particular core allocation coincides with the Vickrey-Clark-Growes (VCG) outcome ([15]) and hence can be implemented in dominant strategies (Corollary 1). On the other hand (Proposition 2), no other core-selecting mechanism is dominant-strategy incentive-compatible. This result gives additional justification for further study of the searcher-optimal point in the core: the value obtained by the validator in this point can be interpreted as the maximal extractable value he can obtain taking into account information rents captured by the searchers.
@@
As a next step, we consider a stochastic model of searcher competition to study the value distribution between the searchers and the validator. Different opportunities (arbitrage or MEV) have a fixed probability of being found by a searcher. The more searchers find an opportunity, the more value can be extracted by the validator and the less value can be extracted by the searchers. We show that if the probability of finding an opportunity is bounded from below by $p>\log(n)/n$, where $n$ is the number of searchers actively searching in the strategy, with high probability the validator captures all value in all core allocations. On the other hand, if the probability of sucessfully searching is very low, $p\in\Theta(1/n)$, then with positive constant probability, searchers will be able to capture all value in the searcher-optimal core allocation.
@@
Next, we empirically validate our stochastic model. We obtain bundle data from the MEV-Share Order Flow Auction (OFA) that allows to study the relation between searcher competition and validator profit. We find that there is significant positive relation between the number of bundles that backrun the same user transaction and the median profit of the validator from opportunities with the same number of backruns.
We extend our model to multiple concurrent proposers (MCP) setting, a recent proposal for block building, which improves economic censorship resistance of the chain, check out [14] and discussion within. In this setting, there are multiple different entities that bring their sets of transactions or bundles, from which the block is formed. We again abstract away from the exact rules how the block is formed from the inputs of proposers. As long as there are at least two searchers, we show that if all but one proposer can include all searcher-relevant transactions in the block, the core of the underlying game is empty. The assumption is reasonable if there are sufficiently many proposers and searcher bundles do not require too much space in the block. This result hinges on the economic instability of multiple concurrent proposers. However, if there is only one searcher, there is a trivial core where all the value goes to the only searcher as the proposers substitute each one of them perfectly.
1.1 Glossary of Terms
In this section, we will list terms that are used throughout the introduction and later in the paper:
  1. 1.
    Miner, Validator, Proposer: A designated party that builds the next block of the blockchain. Initially, the term miner was used, referring to the fact that the next block builder was supposed to find a nonce resulting into a suitable hash of its own block, an activity called ”mining” in the Proof-of-Work context of Bitcoin, [13]. Later, in the Proof-of-Stake context, the role of a block builder was changed to the term validator, also internalizing the functions of validating and voting for the block proposed by the block-proposing entity. In the later evolving block building market of the Ethereum chain, validators usually act as a proposer that signs a block without actually building it, while the parties that actually build blocks are called builders.
  2. 2.
    Arbitrage: Opportunity in which risk-free profits can be made. Usually, arbitrages arise when the same asset is priced differently on multiple markets. In the context of blockchains, the prices between centralized exchange and on-chain exchanges differ, creating arbitrage opportunities in which the arbitrageur sells (buys) over (under) priced asset on chain and buys (sells) under (over) priced asset on the centralized exchange. Different arbitrageurs, however, have different portfolios both on-chain and centralized exchanges, which results into different values they can extract.
  3. 3.
    Searcher: Party specialized in finding arbitrage opportunities. There are searchers that find many different arbitrage types, also searchers that specialize in specific arbitrage types, introducing heterogeneity among them.
  4. 4.
    Frontrunning: Blockchains usually maintain a pool of pending transactions, before they are included and executed in the block(s). If the pool of pending transactions is public, that allows searchers (or proposers/builders) to take advantage by inserting their own transaction in front of a publicly available transaction in the block for execution, leading to a profit.
  5. 5.
    Backrunning: Similar to frontrunning, backrunning refers to inserting the own transaction after a public transaction in the block for execution, leading to a profit for a backrunning party. Backrunning is usually considered a harmless activity, usually fixing the price movement by noise traders, unlike frontrunning.
  6. 6.
    Sandwich attack: Refers to putting the own transactions in front and after a publicly available transaction in the block for execution. This is done by bundling all 3 transactions into one transaction (bundle). Typically, the victim transaction performs an exchange on-chain. The transaction inserted in front increases the price of the asset that the victim transaction buys, and then the following transaction reverts the trade direction. This results in a higher price slippage for a victim transaction sender and gains for an attacker.
  7. 7.
    Liquidation: In many decentralized finance applications, collateral is provided, for example, to borrow some assets. When a collateral asset price goes down, it is usually sold at a disadvantaged price to a depositor, creating a value for a potential buyer. Similarly to arbitrage opportunities, liquidation brings different value to different searchers depending their portfolios.
@@ 1.2 Related Literature
The phenomenon of miner extractable value has first been documented in [5]. Another early contribution on front-running in decentralized finance is [8]. MEV has been documented by a variety of public dashboards and data sets  [^4] For the examples see https://mevboost.pics/,https://libmev.com/. We refer to [19] for the discussion on searchers, builders and validators in MEV extraction, for the Proof-of-Work setting.
Recently, also the magnitude of non-atomic arbitrage has been empirically investigated, where searchers realize one lag of a trade on chain and one on a different domain, see e.g. [10]  [^5] Also check out https://dune.com/flashbots/lol..
The block building market structure that has evolved in Ethereum since the change to proof of stake, has been the topic of several recent contributions: The dashboard https://orderflow.art/ documents empirically the supply chain through which transaction requests land on chain. [1] studies whether MEV and Proof-of-Stake rewards capture leads to centralization, discussing both validator and builder roles in the market formation. [16] argue that in the current market structure searchers and builders have an incentive to vertically integrate. [4] argue that the builder market is prone to centralization. [20] provide descriptive statistics on the level of decentralization on the builder landscape. [3] proposes a dynamic MEV sharing mechanism that the authors argue results into better decentralization and fair allocation. In our paper, we focus on neither the fairness of the value distribution nor long-term centralization threats due to asymmetric value distribution.
@@
The notion of core was first formally defined in [9]. For an overview of results around the core and submodularity, see, e.g. [11]. Similar considerations as in our paper appear in the literature on core-selecting auctions [7, 6].
2 Examples and Model
In this section, we define the game and cooperative game theory terms formally. Before that, we will discuss three illustrative examples to give a motivation and intuition.
2.1 Examples
Suppose that there are three searchers, labeled $s_{1}$, $s_{2}$ and $s_{3}$ and three opportunities, labeled $o_{1}$, $o_{2}$ and $o_{3}$. Consider the following three matrices of values searchers derive for each opportunity. It is assumed that opportunities derive their values independently from each other, and the block size is large enough to accommodate extraction of all opportunities.
S/O $o_{1}$ $o_{2}$ $o_{3}$
$s_{1}$ 11 0 0
$s_{2}$ 0 11 0
$s_{3}$ 0 0 11
In the example one, all searchers find unique opportunities.
S/O $o_{1}$ $o_{2}$ $o_{3}$
$s_{1}$ 0 11 11
$s_{2}$ 11 0 11
$s_{3}$ 11 11 0
In the example two, each opportunity that each searcher finds is found by some other searcher.
S/O $o_{1}$ $o_{2}$ $o_{3}$
$s_{1}$ 11 11 0
$s_{2}$ 0 11 0
$s_{3}$ 0 0 11
In the example three, searcher $s_{1}$ finds one unique opportunity and one opportunity found by the searcher $s_{2}$, which only finds one opportunity, while the searcher $s_{3}$ finds one unique opportunity.
All these examples differ in many aspects, namely, who finds how many opportunities and how unique these opportunities are. One thing that all examples share is that the total value of all opportunities is $3$. It is therefore interesting how this value can be distributed between searchers and the validator that has an exclusive right to propose the next block, taking into account their bargaining power searchers derive from uniquely finding opportunities. In the following section, we define tools to solve this simple problem with independent valuations and a much more general problem with correlated valuations.
2.2 Model
@@
i.e. in case the validator is part of the coalition, the coalitional value is the value of the total welfare maximizing block consisting of transactions from searchers in the coalition, and in case the validator is not part of the coalition, no value can be generated, as the validator is necessary to realize a block. This implicitly assumes a communication mechanism between a validator and searchers, however, the validator does not need to be actively involved into value extraction.
It will be useful subsequently to introduce the short hand notation
@@
if we receive more bundles to build a block that will increase welfare weakly, since we can always discard submitted bundles when building a block.
In all three examples, $\bar{v}(\{s_{1},s_{2},s_{3}\})=3$. In example one, value function is defined as: $\bar{v}(\{s_{i}\})=1$, for any $i$ and $\bar{v}(\{s_{i},s_{j}\})=2$, for any $i\neq j$.
In example two, value function is defined as: $\bar{v}(\{s_{i}\})=2$, for any $i$ and $\bar{v}(\{s_{i},s_{j}\})=3$, for any $i\neq j$.
In example three, value function is defined as: $\bar{v}(\{s_{2}\})=\bar{v}(s_{3})=1$, $\bar{v}(s_{1})=2$, $\bar{v}(\{s_{1},s_{2}\})=\bar{v}(\{s_{2},s_{3}\})=2$, and $\bar{v}(\{s_{1},s3\})=3$.
We make the following assumption on the (collective) value function:
@@
Submodularity states that the value of a block we can build from transactions submitted by searchers in $A$ and $B$, is bounded by subtracting the value of a block we can build from transactions in both $A$ and $B$ from the sum of values we can achieve from building a block with transactions in $A$ and a block with transactions in $B$.
Submodularity requires that there are not-too-strong complementarities between different submitted bundles. We can justify this assumption in two ways: first, it may be that complementarities are not strong and different MEV opportunities provide value that is mostly independent from other opportunities. Second, it may be that the complementarities are already absorbed by searchers, e.g. in the sense that searchers who provide complementary flow have an incentive to integrate their operations and send their flow together to extract more value. It is worth noting that our upper bound on searcher values holds also for non-submodular value functions, but may be loose in that case. Searchers conducting most common arbitrage opportunities described in section 1.1 exhibit no strict complementarities. It can be easily checked that all the examples satisfy the submodularity condition.
@@
Decreasing Marginal Value: Let $A\subseteq B\subseteq\mathcal{S}$ and $a\in A$. Then, the value function satisfies the following inequality:
@@
A direct consequence of submodularity is the following lemma, which will be useful subsequently:
@@
Let ${A}\subseteq{B}\subseteq\mathcal{S}$. Then, for decreasing marginal value functions, the following inequality holds:
@@
We prove the result by induction on $N:=|{B}\setminus{A}|$. For $N=0$ the result holds trivially. Now suppose the result holds for $N\geq 0$ and consider the case $N+1$. Let $j\in B\setminus A$. By induction assumption
@@
We define the allocation of value between all players of the cooperative game.
Definition 1.
@@
Note the abuse of notation, we will use $x_{j}$ instead of $x(j)$ to denote an allocation to a player $j$. Next, we define the core, which is the main solution concept that we will focus on in this paper:
Definition 2.
@@
Inequalities (1) indicate that there is no coalition of players $C$ that get allocated less than the value they can derive themselves, $v(C)$. Intuitively, if such a coalition of players existed, they could profitably deviate from the allocation, leading to instability of the allocation.
@@ 3 Analysis
Our first main result states each searcher can at most capture their marginal contribution to the realized block, and any allocation that gives each searcher at most their marginal contribution is in the core. In particular, this implies that there is a searcher optimal core allocation (giving each searcher exactly their marginal contribution to the realized block) and a validator optimal allocation (giving the validator all realized value).
@@
First, we show that $x_{i}\leq\bar{v}(\mathcal{S})-\bar{v}(\mathcal{S}\setminus\{i\})$ for each $i$. Suppose not. In that case
@@
where the last inequality follows from Lemma 1. Thus, the core inequalities (1) are satisfied for all coalitions $A\cup V$ with $A\subseteq\mathcal{S}$. For coalitions without the validator the core inequalities are trivially satisfied, as all searchers get non-negative value in $x$.
In the example one, the prefix of the core vector $(x_{s_{1}},x_{s_{2}},x_{s_{3}})$ is the product of $0,1]\times[0,1]\times[0,1]$, while the validator is paid $x_{V}=3-x_{s_{1}}-x_{s_{2}}-x_{s_{3}}$.
In the example two, all searchers are getting, i.e., $x_{i}=0$ for any $i\in\{s_{1},s_{2},s_{3}\}$. The validator is paid all the value equal to $3$.
In the example three, the prefix of the core $(x_{s_{1}},x_{s_{2}},x_{s_{3}})$ is the product of $0,1]\times 0\times[0,1]$, while the validator is paid $3-x_{s_{1}}-x_{s_{3}}$.